OpenStudy (anonymous):

Find the slope of a line that contains the points named. P (1,2) Q (3,4)

OpenStudy (jackellyn):

y=mx+b, where m is the slope. m=y2-y1/x2-x1 y2=4 y1=2 x2=3 x1=1 Can you set up the equation?

OpenStudy (precal):